OpenStudy (anonymous):

how to eliminate parameter for x=sint y=csct 0

OpenStudy (aum):

y = csc(t) = 1 / sin(t) = 1 / x

OpenStudy (anonymous):

oh!! lol that was simple!! Thanks! :)

OpenStudy (aum):

You are welcome ;)